# Universal Approximation Theory: The basic theory for large language models

### v1

@@ 3.5 What makes Transformer to be the winner in LLMs?
The Transformer architecture is founded on two pivotal components: FFN and MHA. FFN is comprised of linear operations. In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both Linear and MHA. In this section, we delve into why the Transformer has emerged as a key model in the domain of LLMs.
From a mathematical perspective, both the MLP and the Transformer serve as concrete implementations of the UAT. Through the matrix-vector approach, it is straightforward to demonstrate that convolutional networks also realize the UAT, thus sharing the same foundational mathematical principles. However, the question arises: Why does the Transformer hold a dominant position in LLMs? The answer lies in the ingenious design of the MHA. Unlike traditional linear operations, the parameters of MHA are dynamic, adjusting in response to the input data. While the UAT endows models with a powerful approximation capability, fixing the parameters in linear operations means that the function that an MLP can approximate remains unchanged, as illustrated in Figure 6.a. In contrast, the Transformer can adjust its parameters by $\mathbf{H}_{i}^{\prime}$, based on the input. These parameters, $\mathbf{H}_{i}^{\prime}$, should be regarded as being generated in response to the input, as depicted in Figure 6.b. This adaptability implies that the approximated function can vary with the input. This characteristic also explains the strong generalization capability of Transformer-based neural networks, as they can approximate different functions tailored to the requirements of various tasks, such as those needed for language translation versus mathematical problem-solving.
Figure 6: The approximation ways of the MLP (a) and Transformer (b) networks.
@@ 4 Discussion
Leveranging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and emplore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
4.1 The fundamental reason for the power of Transformers in LLMs
Theoretically, we have established that Transformer networks act as concrete implementations of the UAT, leveraging its potent capability to approximate any Borel measurable function, thus ensuring the high-dimensional function approximation advantage of models based on UAT. While UAT possesses immense function approximation abilities, it inherently lacks the capacity of approximating multiple functions simultaneously. However, language tasks are by nature diverse, potentially requiring the approximation of different functions based on the input. For instance, summarizing, translating, and continuing the same text content, the function space remains entirely the same except for minor variations in prompt words. Without the ability to dynamically approximate functions based on input, but merely fitting based on a general function trend, it would lead to approximating the same function, hence producing identical or similar output. Therefore, LLMs need to distinguish and adapt to these nearly identical functions, thereby dynamically generating response functions based on input. The MHA mechanism in Transformers provides LLMs with the capability to dynamically approximate relevant functions based on input. In particular, MHA mechanism allows for the dynamic adjustment of UAT’s parameters according to the input, thereby laying the theoretical foundation for Transformer-based LLMs to simultaneously tackle multifaceted tasks such as translation, continuation, summarization, code generation, and solving mathematical problems.
@@ 4.2 What enables LLMs to possess ICL capability?
Contextual interaction, as the core capability of LLMs, permeates every phase from training and fine-tuning to prediction. ICL, multi-step reasoning, and instruction following are intuitive manifestations of this contextual interaction. Leveraging their context-sensitive interaction capabilities, LLMs can exhibit behaviors consistent with ICL, multi-step inference, and instruction following, which are tailored based on contextual cues. Furthermore, contemporary LLMs are evolving towards handling longer contexts; however, this advancement is accompanied by a dramatic increase in computational resource demands due to text expansion. A notable solution to this challenge is LongLoRA Chen et al. (2023), which fundamentally also utilizes contextual interaction mechanisms to facilitate training on extended sequences.
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}^{i})^{\prime}$ in Figure 5 reveals this mode of contextual interaction. Since $(\mathbf{W}_{HVO}^{\prime})^{T}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs.
@@ 4.3 What justifies the feasibility of pruning LLMs?
Due to the massive size of parameters in LLMs and the subsequent high demand for computational resources, pruning LLMs is pivotal for their deployment. A legitimate question to ask is why LLMs are amenable to pruning? The rationale lies in the presence of excessively low-weight parameters in certain layers of LLMs. To understand this, we can directly analyze it from the perspective of the formula underlying the UAT:
@@
Figure 7: The LoRA porcess (a), and the operation from the perspective of UAT (b).
@@
4.4 How does the LoRA scheme effectively fine-tune LLMs?
Owing to the substantial computational resources required for training LLMs and their strong generalization capabilities, we believe the focus should be redirected towards efficiently leveraging pre-trained models. Repeatedly training models from scratch incurs considerable computational overhead, rendering the endeavor of repurposing well-trained models in novel tasks both practically valuable and resource-saving. A currently prominent solution addressing this issue is the LoRA methodology.
The LoRA scheme, illustrated in Figure 7.a, adopts a streamlined approach where $T_{i}$ symbolically represents the transformations within the network. Given our prior establishment that these transformations can be cast into matrix-vector operations, from the perspective of the UAT, LoRA fundamentally constitutes a layer-wise fine-tuning tailored to specific task characteristics. This fine-tuning mechanism is depicted in Figures 7.b.1 and Figure 7.b.2, where $\mathbf{W}^{\prime}_{amend}$ signifies the parameters post-refinement, thereby embodying an adaptive calibration of the UAT’s underlying parameters to enhance task-specific performance.
@@ 4.5 Rethinking LLMs
The capabilities of LLMs have become so formidable that their language processing prowess is increasingly approaching human levels, prompting a core question: How do LLMs differ from humans in language processing? Figure 8 elucidates the contrast between the patterns employed by LLMs and humans. Both start with language encoding—humans through a character-based system and LLMs via numerical arrays—essentially, there is little difference at this foundational level. Given the polysemy of words, context determination is crucial: humans understand context through the activation and transmission of neurons in the human brain, while LLMs use patterns of parameter activation to approximate corresponding functions. Here, the activation of network parameters parallels the neural electrical signalling in the brain, and the approximation of functions corresponds to human comprehension of semantics. From this perspective, the difference in language processing between humans and LLMs seems minimal.
We posit that the main differences lie in human autonomy and multimodal perception. Human autonomy enables self-directed learning, allowing knowledge to be validated within real-world contexts and personal experiences, whereas LLMs are confined to function fitting based on the majority of corpora in their datasets, with the goal of optimizing loss minimization. Human multimodality encompasses interpreting the world through senses such as vision, hearing, taste, and touch, infusing vocabulary with rich meanings, while LLMs are inherantly adept to to single-modality numerical representations.
Another significant issue in LLMs pertains to the segmentation of word embeddings within the MHA, as shown in Figure 4.a. Does this fragmentation still capture the full meaning of words comprehensively? If there are eight attention heads, does that mean a word has eight different encodings, suggesting a form of modality diversity, which then interacts through MHA to extract contextual insights? Furthermore, by considering LLMs as implementations of UAT, can we directly compute the weights necessary for function approximation from the data, thereby significantly reducing training overhead?
@@ 5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules, which enable the adaptation to and approximation of diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s innate capacity to learn from context.
@@
Limitation
Our work primarily delves into explaining a select few characteristics (ICL, Instruction following, and multi-step reasoning) and techniques (such as LoRA and Pruning) that we perceive as pivotal within current LLMs. While numerous advancements have been made to Transformer-based LLMs, their fundamental underpinnings largely align with the principles outlined herein. Nevertheless, due to space constraints, this paper does not aspire to exhaustively address every issue or attribute present across all LLMs. Instead, it focuses on elucidating those deemed most significant, thereby offering a concentrated insight into the core of LLMs. It is important to acknowledge that our scope, though carefully curated, is by no means exhaustive, and further exploration is warranted to gain a comprehensive understanding of the extensive landscape of LLMs enhancements and challenges.
@@ Appendix A The Property of Diamond
Appendix B The Matrix-Vector Form of Linear and MHA
In this section, we will present the transformation process Linear and MHA into their corresponding matrix-vector forms.
B.1 The Matrix-Vector Form of Linear
Eq. ((15)) gives the transfomation method for Linear, it can be deduced that if the input data matrix has dimensions $(N,M)$ and the parameter matrix has dimensions $(N,N)$, the sparsity of the parameter matrix corresponding to the Linear operation is $1/M$.
$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&\cdots&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{n,1}&w_{n,2}&\cdots&w_{n,n}\end{array}\right)\left(\begin{array}[]{cccc}x_{1,1}&x_{1,2}&\cdots&x_{1,m}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ x_{n,1}&x_{n,2}&\cdots&x_{n,m}\end{array}\right)=\left(\begin{array}[]{cccc}y_{1,1}&y_{1,2}&\cdots&y_{1,m}\\ y_{2,1}&y_{2,2}&\cdots&y_{2,m}\\ \vdots&\vdots&\vdots&\vdots\\ y_{n,1}&y_{n,2}&\cdots&y_{n,m}\end{array}\right)$ (15)
$\rightarrow\left(\begin{array}[]{ccc}w_{1,1}&0&\cdots\\ 0&w_{1,1}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ w_{1,2}&0&\cdots\\ 0&w_{1,2}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\\ \vdots&\vdots&\cdots\\ w_{1,n}&0&\cdots\\ 0&w_{1,n}&\cdots\\ \vdots&\vdots&\cdots\\ 0&0&\cdots\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1,1}\\ x_{1,2}\\ \vdots\\ x_{1,m}\\ x_{2,1}\\ x_{2,2}\\ \vdots\\ x_{2,m}\\ \vdots\\ x_{n,1}\\ x_{n,2}\\ \vdots\\ x_{n,m}\end{array}\right)=\left(\begin{array}[]{c}y_{1,1}\\ y_{1,2}\\ \vdots\\ y_{1,m}\\ y_{2,1}\\ y_{2,2}\\ \vdots\\ y_{2,m}\\ \vdots\\ y_{n,1}\\ y_{n,2}\\ \vdots\\ y_{n,m}\end{array}\right)$
B.2 The Matrix-Vector Form of MHA
@@
Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.

### v3

@@ 3.5 The UAT Format of Multi-Layer Transformer
The Transformer architecture is founded on two pivotal components: FFN and MHA. In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both FFN and MHA. In this section, we delve into why multi-layer Transformer is the implementation of UAT.
Based on Equations 6 and 10, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer is a concrete realization of the UAT, which can be written as (see Appendix C for more details):
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input. This enhances the model’s ability to dynamically adjust functions based on input. In the MHA mechanism, the parameters change dynamically with the input. Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
@@ 4 Discussion
Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
4.1 What Makes Transformer So Powerful in LLMs?
Theoretically, we have established that Transformer networks are concrete implementations of the UAT, enabling them to approximate any Borel measurable function. While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously. However, language tasks are inherently diverse, often requiring the approximation of different functions based on the input. For instance, when summarizing, translating, or continuing the same text, the input functions are nearly identical, with only minor variations in the prompt. Without the ability to dynamically approximate functions based on the input, simply fitting a general function trend based on input will result in identical or similar outputs.
To address this, LLMs must distinguish and adapt to these nearly identical functions, dynamically generating response functions based on the input. The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input. Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms, as illustrated in Figure 6. This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
Figure 6: This diagram vividly demonstrates the output differences between an MLP and a Transformer when given the same input content (represented by the blue line in a) with only slight variations (represented by the black line in a). Since the input content is nearly identical and the MLP’s parameters are fixed, the MLP tends to produce the same output (b), influenced by the overall input function trend. However, the Transformer, with its ability to dynamically adapt to the input, generates different outputs based on the input variations. In the context of LLMs, the blue line in a can be understood as a consistent piece of text, while the black line represents different prompts, such as translation, continuation, or summarization.
Furthermore, the MHA mechanism can capture global information (as shown in Figure 5, where each element in the output $\mathring{\mathbf{H}}^{\prime}$ contains global information), which helps generate content that is consistent with the context. This is crucial for language understanding that requires extensive contextual information. In contrast, 1D convolution employs a sliding convolution learning strategy, where the learned content is somewhat influenced by the size of the convolution kernel, leading to a more fragmented learning approach (for more information, please refer to Appendix D).
@@ 4.2 What Enables LLMs to Possess ICL Capability?
Contextual interaction, as the core capability of LLMs, permeates every phase from training and fine-tuning to prediction. ICL, multi-step reasoning, and instruction following are intuitive manifestations of this contextual interaction. Leveraging their context-sensitive interaction capabilities, LLMs can exhibit behaviors consistent with ICL, multi-step inference, and instruction following, which are tailored based on contextual cues.
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ in Figure 5 reveals this mode of contextual interaction. Since $\mathbf{W}^{\prime}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs. (See Appendix D for more details)
@@ 4.3 What Justifies the Feasibility of Pruning LLMs?
Due to the massive size of parameters in LLMs and the subsequent high demand for computational resources, pruning LLMs is pivotal for their deployment. A legitimate question to ask is why LLMs are amenable to pruning. The rationale lies in the presence of excessively low-weight parameters in certain layers of LLMs. To understand this, we can directly analyze it from the perspective of the formula underlying the UAT:
@@
4.4 How Does the LoRA Scheme Effectively Fine-tune LLMs?
Given the substantial computational resources required to train LLMs and their powerful generalization abilities, we believe that more efficient use of pre-trained models is essential. Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient. A prominent solution to this challenge is the LoRA (Hu et al. 2021), which can be expressed as follows:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}$ (16)
According to Eq. 11, we use LoRA to fine-tune the Linear operation in FFN can be written as:
$\mathbf{x}_{i+1}$ $=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}(\mathbf{W}^{\prime}_{j,3}+\Delta\mathbf{W}_{j+1,3})$ (17)
$\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
From Eq. 17, it can be seen that LoRA essentially fine-tunes the UAT parameters layer by layer for a specific task.
@@ 4.5 Rethinking LLMs
The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing? Figure 7 illustrates the comparison between the language processing processes of LLMs and humans. Both start with language encoding—humans encode language through a character-based system, while LLMs use numerical arrays. At this level, there is almost no difference. Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using UAT. Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans. From this perspective, the differences between humans and LLMs in language processing seem minimal.
So how do we explain human understanding, analysis of language, and memory retrieval? Are these also capabilities of LLMs? First, we can consider the human brain as a combination of one or more UAT models, which are randomly initialized at birth. What supports such an assumption? The original neural network, the perceptron (Rosenblatt 1963), was designed based on the human neuron. Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons. However, through our derivation, the Transformer can also be understood as an implementation of UAT, and the perceptron can be seen as a layer of UAT. The human brain can be considered as a cluster of multi-layer perception. Learning language in a social environment is akin to training the UATs in the brain. Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs. Understanding and analysis: When we respond to a question posed by someone, it can be seen as a simple example of understanding and analysis, which both humans and LLMs can accomplish. Humans may guess the answer to unfamiliar questions based on their experiences, which can be either correct or incorrect, and LLMs exhibit similar behavior. Memory retrieval: This is essentially fitting specific results based on particular words, as there is no actual database in the brain. For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface. This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E). However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions.
Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face. We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple UAT models to produce reasonable results). Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs. LLMs, on the other hand, are limited to function approximation within the corpora data.
@@ 5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s MHA to learn from context.
@@ Appendix A The Property of Diamond
Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 2: The process of Transformer.
Appendix B The Matrix-Vector Form of MHA
In this section, we will present the transformation process of MHA into their corresponding matrix-vector forms.